When theTitanic sank in 1912,it weighed 52,310 tons.Its weight years after sinking satisfies the equation

weight = Ae-0.0023t a) How much does the Titanic weigh in 2021?

     b) How many years from now will theTitanic weigh10%of its original weight – 5231 tons?

  


1
Expert's answer
2021-06-17T11:32:47-0400

a) The Titanic weight in 2021 was:

"weight = 52,310*e^{-0.0023*(2021 - 1912)} = 40,710.56."

b) The number of years from now for theTitanic to weigh 10% of its original weight – 5231 tons is:

"52,310*e^{-0.0023*(109 + t)} = 5,231,"

"e^{-0.0023(109 + t)} = 0.1,"

t = 892.1 years.